Please help!!
I need answers to the following questions, I really need to past this test.
1) In solving the equation (x + 1)(x – 2) = 54, Eric stated that the solution would be
x + 1 = 54 => x = 53 or (x – 2) = 54 => x = 56
However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.
2)If a stone is tossed from the top of a 190 meter building, the height of the stone as a function of time is given by h(t) = -9.8t^2 -10t +190, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer.
3) Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
8x^2 + 2x + 4 = 0
A) Two different irrational solutions
B) Two different imaginary solutions
C)Exactly one rational solution
D)Two different rational solutions
4) Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.
18x + 6y = 78
12x + 54y = -48
5)Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
14x+18y =-54
-4x-14y = 42
Please I need the answers to these questions ASAP like yesterday. Been trying all day to figure out the answers to the above 5 questions, thank you

Answer by sudhanshu_kmr(1152)


Such equations are called quadratic equations. Use factorization method or
quadratic formula to solve such equations.
(x + 1)(x – 2) = 54
=> x^2 -2x + x -2 = 54
=> x^2 -x -56 =0
=> x^2 -8x +7x -56 =0
=> x(x-8) + 7(x -8)= 0
=>(x+7)(x-8) = 0
x= -7, or 8 .